# ]organization=LASIGE, Faculdade de Ciências, Universidade de Lisboa , country=Portugal

### v2

@@ 1 Introduction
This paper presents an empirical study that follows a comparative approach to answer the following research question: are LLM chatbots competitive with state-of-the-art specialized models for detecting OSINT CTI and extracting pertinent information? To produce the empirical results and evaluate the performance of chatbots, we use a publicly-available annotated dataset [10, 11] collected from Twitter, a known reliable OSINT CTI source [12]. This dataset provides an accurate testing environment for challenging chatbots in a real-world CTI scenario.
@@ 2.4 OSINT for cybersecurity
Although there are many ML techniques for binary classification and NER [8, 9], to the best of our knowledge, the architecture of Dionísio et al. [11, 41] is recognized [42] as the state-of-the-art for OSINT-based CTI extraction on Twitter data, reaching almost 95% for different quality metrics. This work also shows superior performance when compared with classifiers based on SVMs and Multi-Layer Perceptrons (MLPs), and several alternative NER approaches. Due to this, we consider Dionisio et al. [41] as the reference specialized model in our comparison.
@@ 4 Chatbots evaluation
We divided this section into three topics relevant to the evaluation: dataset, methodology, and evaluation criteria. First, we introduce the Twitter dataset, which is the foundational source for generating subsequent prompts. The methodology subsection outlines the techniques and strategies used to assess and compare chatbot performances. Finally, in the evaluation criteria subsection, we explore the metrics and standards by which the chatbots are assessed.
@@ 6.2 Evaluation of NER
Approach Number of Questions
Entity
$\text{F}_{1}$ score Execution Time
ESP 11074
Organization
0.36 4h 02m
ESP 11074
Version
0.43 4h 23m
GLP 11074
All entities
0.10 3h 09m
@@ Acknowledgement
timestamp keywords
original tweet
pre-processed tweet relevance entities
2018-07-24 01:00:46+00:00 oracle
RT Oracle: Learn to use and understand #Oracle’s Internet Intelligence Map https://t.co/l06Nyf1FFF Dyn https://t.co/uzozFKwm97
rt oracle learn to use and understand oracle s internet intelligence map dyn 0 -
2016-12-09 19:19:38+00:00 internet explorer
threatmeter: [dos] - Microsoft Internet Explorer 9 MSHTML - CDisp Node::Insert Sibling Node Use-After-Free (MS13-0… https://t.co/gLvEwpDL9v
threatmeter dos microsoft internet explorer 9 mshtml cdisp node::insert sibling node use-after-free ms13-0 1 O O B-ORG B-PRO I-PRO B-VER O O O O O B-VUL B-ID

### v3

@@ 1 Introduction
This paper presents an empirical study that uses inductive reasoning and a comparative methodology to answer the following research question: Are LLM chatbots competitive with state-of-the-art specialized models for detecting OSINT CTI and extracting pertinent information? To produce the empirical results and inductively assess the performance of chatbots, we use a publicly available annotated dataset [10, 11] collected from Twitter, a known reliable OSINT CTI source [12]. This dataset provides an accurate testing environment for challenging chatbots in a real-world CTI scenario. Then, a comparative analysis of previous work and this paper’s results effectively answers the research question, allowing the discussion of the conclusion’s implications and outlining possibilities for further research and improvements.
@@ 2.4 OSINT for cybersecurity
Although there are many ML techniques for binary classification and NER [8, 9], to the best of our knowledge, the architecture of Dionísio et al. [11, 41] is recognized [42] as the state-of-the-art for OSINT-based CTI extraction on Twitter data, reaching almost 95% for different quality metrics. This work also shows superior performance when compared with classifiers based on SVMs and Multi-Layer Perceptrons (MLPs), and several alternative NER approaches. Due to this, we consider Dionisio et al. [41] as the reference specialized model in our comparative analysis.
@@ 4 Chatbots evaluation
We divided this section into three topics relevant to the evaluation: dataset, experimental methodology, and experimental results evaluation criteria. First, we introduce the Twitter dataset, which is the foundational source for generating subsequent prompts. The methodology subsection outlines the techniques and strategies used to assess and compare chatbot performances. Finally, in the evaluation criteria subsection, we explore the metrics and standards by which the chatbots are assessed.
@@ 6.2 Evaluation of NER
{tblr}
colspec=lccc, roweven = c, row1 = Mercury, row3 = c, cell11 = c, cell12 = c, cell14 = c, cell15 = c, cell16 = c, hlines, Approach & Number of Questions Entity $\text{F}_{1}$ score Execution Time
ESP 11074 Organization 0.36 4h 02m
ESP 11074 Version 0.43 4h 23m
GLP 11074 All entities 0.10 3h 09m
@@ Acknowledgement
{tblr}
colspec = X[-1,l] l X[2,l] X[2,l] c X[l], row1 = Alto,c, cell22 = c, cell25 = c, cell26 = c, cell32 = c, cell35 = c, hline1-4 = -, timestamp & keywords original tweet pre-processed tweet relevance entities
2018-07-24 01:00:46+00:00 oracle RT Oracle: Learn to use and understand #Oracle’s Internet Intelligence Map https://t.co/l06Nyf1FFF Dyn https://t.co/uzozFKwm97 rt oracle learn to use and understand oracle s internet intelligence map dyn 0 -
2016-12-09 19:19:38+00:00 internet explorer threatmeter: [dos] - Microsoft Internet Explorer 9 MSHTML - CDisp Node::Insert Sibling Node Use-After-Free (MS13-0… https://t.co/gLvEwpDL9v threatmeter dos microsoft internet explorer 9 mshtml cdisp node::insert sibling node use-after-free ms13-0 1 O O B-ORG B-PRO I-PRO B-VER O O O O O B-VUL B-ID